Given the following SOP equation.

F = Σ (2,4,5,6,10,12,13,14)

Fill-out the Truth Table and write the unsimplified SOP equation. FEQ FCircuit 11 11

FEQ=

2- Draw the logic circuit for the unsimplified SOP equation.

3-

Using Boolean algebra, simplify the SOP equation (FEQ) and draw the logic circuit for the simplified SOP equation.

F =
Answer #1

1) Truth F- table - - - - - --- e coooooo > - - - - c o º o - - - - 0 0 0 0 0 - -oo-oo- -oo--oos -0-0-0-0-0---00 o - -- -ooo

EN9 un simplified sop equation is FSO = Ā B C D + Ā B C D + ĀB C D + Ā B C Õ+ A B C D + A B E Ď + A B C D + A B C o 2) LogicÕ+ A B C r + F = Ā B C ŏ + Ā B C D + A B C Ā BET + Ā B C D + A B C D + A B C D we have A+ A=1 A B = Ā B + AB B A B = AB + AB